# When to use Graphs in RAG: A Comprehensive Analysis for Graph Retrieval-Augmented Generation

### v1

@@ Abstract
Graph retrieval-augmented generation (GraphRAG) has emerged as a powerful paradigm for enhancing large language models (LLMs) with external knowledge. It leverages graphs to model the hierarchical structure between specific concepts, enabling more coherent and effective knowledge retrieval for accurate reasoning. Despite its conceptual promise, recent studies report that GraphRAG frequently underperforms vanilla RAG on many real-world tasks. This raises a critical question: Is GraphRAG really effective, and in which scenarios do graph structures provide measurable benefits for RAG systems? To address this, we propose GraphRAG-Bench, a comprehensive benchmark designed to evaluate GraphRAG models on both hierarchical knowledge retrieval and deep contextual reasoning. GraphRAG-Bench features a comprehensive dataset with tasks of increasing difficulty, covering fact retrieval, complex reasoning, contextual summarization, and creative generation, and a systematic evaluation across the entire pipeline, from graph construction and knowledge retrieval to final generation. Leveraging this novel benchmark, we systematically investigate the conditions when GraphRAG surpasses traditional RAG and the underlying reasons for its success, offering guidelines for its practical application. All related resources and analyses are collected for the community at https://github.com/GraphRAG-Bench/GraphRAG-Benchmark.
@@ 1 Introduction
Large language models (LLMs), like Claude anthropic2024claude and GPT openai2023gpt4 series, have surprised the world with their remarkable capabilities in many real-world tasks, like linguistic comprehension brown2020language , question answering khashabi2020unifiedqa , mathematical reasoning hong2025benchmarking , and content generation chowdhery2023palm ; hong2024next ; zhang2024structure . Despite the success, LLMs are always criticized for their inability to handle knowledge-intensive tasks and the tendency to generate hallucinations huang2023survey , especially when faced with questions requiring specialized expertise LLM4EA ; he2024chineseqa ; tan2024chineseqa . Retrieval-augmented generation (RAG) gao2023retrieval ; lewis2020retrieval has recently offered a promising approach to adapt LLMs for specific or private domains. Rather than retraining LLMs to incorporate new knowledge and updates feng2025geoedit ; fang2024alphaedit ; jiang2025anyedit ; wang2024knowledge , RAG enhances these models by leveraging external knowledge from text corpora. This approach enables LLMs to generate responses by leveraging not only their parametric knowledge but also real-time retrieved domain-specific information, thereby providing more accurate and reliable answers chen2024multi ; li2024structrag .
@@
However, traditional RAG systems often face critical challenges when dealing with large-scale, unstructured domain corpora edge2024local ; peng2024graph . The textual documents in this corpus, collected from different sources, like research papers, textbooks and technical reports, often vary widely in accuracy and completeness guo2025empowering . The information retrieved by RAG systems can be extensive, complex, and lack clear organization, since domain knowledge is typically scattered across multiple documents without clear hierarchical relationships between different concepts sun2024thinkongraph ; zhang2024knowgpt ; ma2024think-on-graph-2.0 . Although RAG systems borgeaud2022improving ; izacard2023atlas ; jiang2023active attempt to manage this complexity by dividing documents into smaller chunks for effective indexing, this approach inadvertently sacrifices crucial contextual information, significantly compromising retrieval accuracy and contextual comprehension for complex reasoning han2024retrieval ; zhang2025survey ; difflogic .
@@
To address this, graph retrieval-augmented generation (GraphRAG) zhang2025survey ; peng2024graph ; procko2024graph has recently emerged as a powerful paradigm that leverages external structured graphs to improve LLMs’ capability on contextual comprehension han2024retrieval ; zhang2025survey . Early efforts, like Microsoft GraphRAG edge2024local and its variant LazyGraphRAG LazyGraphRAG , employ hierarchical community-based search and combine local/global querying for comprehensive responses. Building on this, LightRAG guo2024lightrag improves scalability through dual-level retrieval and graph-enhanced indexing, while GRAG hu2024grag introduces a soft pruning technique to mitigate the impact of irrelevant entities in retrieved subgraphs and employs graph-aware prompt tuning to help LLMs interpret topological structure. Further extending these capabilities, StructRAG li2024structrag tailors data structures to specific tasks by dynamically selecting optimal graph schemas, while KAG liang2024kag constructs domain expert knowledge using conceptual semantic reasoning and human-annotated schemas, which significantly reduces noise present in OpenIE systems. These strategies used in GraphRAG models significantly improve retrieval precision and contextual depth, enabling LLMs to address complex, multi-hop queries more effectively.
@@
Despite its conceptual promise, recent studies han2025rag ; zhou2025depth report that GraphRAG models frequently underperform traditional RAG approaches on many real-world tasks. Specifically, the previous study han2025rag demonstrates that GraphRAG achieves 13.4% lower accuracy on the Natural Question compared to vanilla RAG, with particularly poor performance on time-sensitive queries (e.g., 16.6% accuracy drop for questions requiring real-time knowledge updates). While graph-based retrieval improves reasoning depth by only 4.5% on HotpotQA’s multi-hop questions, it introduces 2.3 × higher latency on average zhou2025depth . These inconsistencies between conceptual potential and practical efficacy raise critical questions: Is GraphRAG really effective, and in which scenarios do graph structures provide measurable benefits for RAG systems?
@@
It is crucial to identify the factors that are currently limiting GraphRAG’s real-world performance. However, quantitatively and fairly assessing the role of graph structures in RAG systems is challenging. Current benchmarks, including HotpotQA yang2018hotpotqa , MultiHopRAG tang2024multihop and UltraDomain qian2024memorag , fail to adequately evaluate the effectiveness of graph structures in RAG systems due to fundamental limitations in both their problem design and corpus composition. ❶ First, existing benchmarks lack granular differentiation in task complexity. Existing benchmarks overemphasize retrieval difficulty, locating scattered facts from corpora, while neglecting reasoning complexity, which involves synthesizing interconnected facts into contextually grounded solutions. As shown in Figure 3, they predominantly focus on narrow task categories, such as simple fact retrieval or linear multi-hop reasoning, without systematically capturing the spectrum of challenges encountered in real-world scenarios tang2024multihop . For instance, a typical multi-hop question in existing benchmarks might ask, “Who founded Company Kjaer Weis, and in which city was this person born¿‘ This requires only the extraction of several discrete facts and cannot extend to complex scenarios requiring hierarchical reasoning and contextual synthesis. ❷ Second, corpora in existing RAG benchmarks suffer from inconsistent quality and low information density. Many datasets are built on generic sources like Wikipedia or news articles, which lack domain-specific knowledge or explicit logical connections. While some work, like UltraDomain qian2024memorag , has tried to extract domain-specific corpora from textbooks, they often fail to encode implicit hierarchies for real-world applications. This makes it impossible to assess GraphRAG’s core strengths in leveraging domain hierarchies. For example, a corpus with poorly defined conceptual hierarchies or loosely connected entities cannot meaningfully test whether graph-aware retrieval mechanisms improve multi-hop reasoning or preserve contextual coherence during knowledge acquisition. Additionally, the absence of corpora with varying information densities, ranging from tightly structured domain knowledge to loosely organized real-world documents, further restricts the evaluation of graph structures’ scalability and adaptability.
@@
To bridge this gap, we propose GraphRAG-Bench, a comprehensive benchmark designed to evaluate GraphRAG models on deep reasoning. GraphRAG-Bench features ❶ comprehensive corpora with different information density, including tightly structured domain knowledge and loosely organized texts, and ❷ tasks of increasing difficulty, covering fact retrieval, multi-hop reasoning, contextual summarization, and creative generation, and ❸ systematic evaluation across the entire pipeline, from graph construction and knowledge retrieval to final generation. Leveraging this novel benchmark, we systematically investigate the conditions when GraphRAG surpasses traditional RAG systems and the underlying reasons for its success, offering guidelines for its practical application.
@@ 2 Preliminary Study
We carefully compare GraphRAG’s pipeline with traditional RAG’s and summarize their characteristics in Figure 1. Generally, RAG retrieves contextually relevant data from a corpus during inference, enabling real-time, domain-specific responses without model retraining. While efficient, its reliance on direct semantic similarity may overlook the broader contextual web of relationships, hierarchies, or implicit logic that binds concepts together. GraphRAG addresses this limitation by expanding the retrieval framework beyond semantic relevance. It structures background knowledge as a graph, where nodes represent entities, events, or themes, and edges define their logical, causal, or associative connections. When processing a query, GraphRAG retrieves not only directly related nodes but also traverses the graph to capture interconnected subgraphs, uncovering latent patterns such as thematic evolution, indirect dependencies, or multi-step reasoning chains. This approach enables the model to synthesize insights from dispersed data points, making it particularly good at tasks demanding complex logical inference. For instance, while RAG might retrieve isolated facts about a topic, GraphRAG could identify related events, causal chains, or thematic clusters, thereby enabling more coherent and comprehensive responses. To sum up, the primary distinction between these two paradigms lies in their handling of contextual depth. RAG excels in scenarios requiring rapid access to discrete information, while GraphRAG emphasizes deep contextual analysis for tasks requiring nuanced understanding of interconnected data. More detailed analysis is included in Appendix G.
@@ 2.2 Current RAG Benchmarks
Existing benchmarks, such as HotpotQA yang2018hotpotqa , MultiHopRAG tang2024multihop and UltraDomain qian2024memorag , were primarily designed to evaluate traditional text-centric RAG frameworks. While these benchmarks have advanced the field, they exhibit critical limitations when applied to assessing GraphRAG.
First, existing benchmarks narrowly focus on testing retrieval difficulty, the ability to locate scattered information from the corpus, while neglecting the equally critical challenge of reasoning difficulty, which involves integrating interconnected concepts/facts by capturing the latent logic. While these benchmarks include “multi-hop“ questions to test a model’s ability in complex reasoning, they do not reflect real-world scenarios demanding complex logical synthesis. For instance, a typical multi-hop question in existing benchmarks might ask, “Who founded Company Kjaer Weis, and in which city was this person born¿‘ This requires only the extraction of several discrete facts from the corpus. However, real-world problems, such as explaining why Company Kjaer Weis failed in a specific market, demand synthesizing financial reports, competitor analyses, consumer trends, and regulatory changes into a coherent narrative. GraphRAG’s strength lies in mapping these interdependencies (e.g., “Market entry timing → supply chain disruptions → regulatory fines → brand erosion“) through graph traversals. However, current benchmarks lack tasks that explicitly require such synthesis, reducing “multi-hop“ queries to sequential fact retrieval within narrow contexts, failing to evaluate how models infer domain-specific hierarchies.
@@
Figure 2: Average number of entities across benchmarks.
@@
Second, the corpora used in existing benchmarks suffer from inconsistent quality and low information density. Most datasets are built on generic sources like Wikipedia or news articles, which lack structured domain-specific knowledge or explicit logical connections. A corpus with poorly defined conceptual hierarchies or loosely connected entities cannot meaningfully test whether graph-aware retrieval mechanisms improve multi-hop reasoning or preserve contextual coherence during knowledge integration. While some work, like UltraDomain, has tried to construct domain-specific corpora using textbooks, they often fail to encode implicit hierarchies for real-world applications. As shown in Table LABEL:tab:pre_sta and Appendix E, domain concepts and their hierarchical dependency appear sparsely in the corpus. This sparsity falls far below the threshold of multi-hop reasoning, which makes it impossible to assess GraphRAG’s core strengths in leveraging domain hierarchies. Additionally, the absence of corpora with varying information densities, ranging from tightly structured domain knowledge to loosely organized real-world documents, further restricts the evaluation of graph structures’ scalability.
Third, current benchmarks fall short in evaluating GraphRAG since their evaluation metrics focus solely on the final outputs, answer accuracy or fluency, while treating GraphRAG’s internal processes (graph construction, retrieval, and generation) as black boxes. Such evaluations can hardly measure how graph structures contribute to the retrieval and reasoning processes. To truly assess GraphRAG, a more holistic evaluation is necessary, encompassing the entire pipeline. This includes examining the efficiency of graph construction and the quality of the resulting graph, the structure and relevance of the knowledge retrieved via the graph, and finally, the faithfulness of the generated answer to this graph-derived context. Such a comprehensive view is essential to understand the actual impact and benefits of graph structures within RAG systems.
@@ 3 GraphRAG-Bench
In this section, we present GraphRAG-Bench, a novel benchmark specifically designed to assess GraphRAG systems through comprehensive task hierarchies and structured knowledge integration. Specifically, GraphRAG-Bench consists a comprehensive dataset with (i) tasks of increasing difficulty, covering fact retrieval, multi-hop reasoning, contextual summarization, and creative generation, and (ii) real-world corpora with different information density, and (iii) a systematic evaluation across the entire pipeline, from graph construction and knowledge retrieval to final generation.

### v2

@@ Abstract
Graph retrieval-augmented generation (GraphRAG) has emerged as a powerful paradigm for enhancing large language models (LLMs) with external knowledge. It leverages graphs to model the hierarchical structure between specific concepts, enabling more coherent and effective knowledge retrieval for accurate reasoning. Despite its conceptual promise, recent studies report that GraphRAG frequently underperforms vanilla RAG on many real-world tasks. This raises a critical question: Is GraphRAG really effective, and in which scenarios do graph structures provide measurable benefits for RAG systems? To address this, we propose GraphRAG-Bench, a comprehensive benchmark designed to evaluate GraphRAG models on both hierarchical knowledge retrieval and deep contextual reasoning. GraphRAG-Bench features a comprehensive dataset with tasks of increasing difficulty, covering fact retrieval, complex reasoning, contextual summarize, and creative generation, and a systematic evaluation across the entire pipeline, from graph construction and knowledge retrieval to final generation. Leveraging this novel benchmark, we systematically investigate the conditions when GraphRAG surpasses traditional RAG and the underlying reasons for its success, offering guidelines for its practical application. All related resources and analysis are collected for the community at https://github.com/GraphRAG-Bench/GraphRAG-Benchmark.
@@ 1 Introduction
Large language models (LLMs), like Claude (Anthropic, 2024) and GPT (OpenAI, 2023) series, have surprised the world with their remarkable capabilities in many real-world tasks, like linguistic comprehension (Brown et al., 2020), question answering (Khashabi et al., 2020), mathematical reasoning (Hong et al., 2025), and content generation (Chowdhery et al., 2023; Hong et al., 2024; Zhang et al., 2024a). Despite the success, LLMs are always criticized for their inability to handle knowledge-intensive tasks and the tendency to generate hallucinations (Huang et al., 2023), especially when faced with questions requiring specialized expertise (Chen et al., 2024b; He et al., 2024; Tan et al., 2024). Retrieval-augmented generation (RAG) (Gao et al., 2023; Lewis et al., 2020) has recently offered a promising approach to adapt LLMs for specific or private domains. Rather than retraining LLMs to incorporate new knowledge and updates (Feng et al., 2025; Fang et al., 2025; Jiang et al., 2025; Wang et al., 2024b), RAG enhances these models by leveraging external knowledge from text corpora. This approach enables LLMs to generate responses by leveraging not only their parametric knowledge but also real-time retrieved domain-specific information, thereby providing more accurate and reliable answers (Chen et al., 2024a; Li et al., 2024).
@@
However, traditional RAG systems often face critical challenges when dealing with large-scale, unstructured domain corpora (Edge et al., 2024; Peng et al., 2024). The textual documents in this corpus, collected from different sources, like research papers, textbooks and technical reports, often vary widely in accuracy and completeness (Guo et al., 2025). The information retrieved by RAG systems can be extensive, complex, and lack clear organization, since domain knowledge is typically scattered across multiple documents without clear hierarchical relationships between different concepts (Sun et al., 2024; Zhang et al., 2024b; Ma et al., 2024). Although RAG systems (Borgeaud et al., 2022; Izacard et al., 2023; Jiang et al., 2023) attempt to manage this complexity by dividing documents into smaller chunks for effective indexing, this approach inadvertently sacrifices crucial contextual information, significantly compromising retrieval accuracy and contextual comprehension for complex reasoning (Han et al., 2024; Zhang et al., 2025; Shengyuan et al., 2024).
@@
To address this, graph retrieval-augmented generation (GraphRAG) (Zhang et al., 2025; Peng et al., 2024; Procko & Ochoa, 2024) has recently emerged as a powerful paradigm that leverages external structured graphs to improve LLMs’ capability on contextual comprehension (Han et al., 2024; Zhang et al., 2025). Early efforts, like Microsoft GraphRAG (Edge et al., 2024) and its variant LazyGraphRAG (Darren Edge, 2024), employ hierarchical community-based search and combine local/global querying for comprehensive responses. Building on this, LightRAG (Guo et al., 2024) improves scalability through dual-level retrieval and graph-enhanced indexing, while GRAG (Hu et al., 2024) introduces a soft pruning technique to mitigate the impact of irrelevant entities in retrieved subgraphs and employs graph-aware prompt tuning to help LLMs interpret topological structure. Further extending these capabilities, StructRAG (Li et al., 2024) tailors data structures to specific tasks by dynamically selecting optimal graph schemas, while KAG (Liang et al., 2024) constructs domain expert knowledge using conceptual semantic reasoning and human-annotated schemas, which significantly reduces noise present in OpenIE systems. These strategies used in GraphRAG models significantly improve retrieval precision and contextual depth, enabling LLMs to address complex, multi-hop queries more effectively.
@@
Despite its conceptual promise, recent studies (Han et al., 2025; Zhou et al., 2025) report that GraphRAG models frequently underperform traditional RAG approaches on many real-world tasks. Specifically, the previous study (Han et al., 2025) demonstrates that GraphRAG achieves 13.4% lower accuracy on Natural Question compared to vanilla RAG, with particularly poor performance on time-sensitive queries (e.g., 16.6% accuracy drop for questions requiring real-time knowledge updates). While graph retrieval improves reasoning depth by 4.5% on HotpotQA’s multi-hop questions, it introduces 2.3 × higher latency on average (Zhou et al., 2025). These inconsistencies between conceptual potential and practical efficacy raise critical questions: Is GraphRAG really effective, and in which scenarios do graph structures provide measurable benefits for RAG systems?
@@
It is crucial to identify the factors that are currently limiting GraphRAG’s real-world performance. However, quantitatively and fairly assessing the role of graph structures in RAG systems is challenging. Current benchmarks, including HotpotQA (Yang et al., 2018), MultiHopRAG (Tang & Yang, 2024) and UltraDomain (Qian et al., 2024), fail to adequately evaluate the effectiveness of graph structures in RAG systems due to fundamental limitations in both their problem design and corpus composition. ❶ First, existing benchmarks lack granular differentiation in task complexity. Existing benchmarks overemphasize retrieval difficulty, locating scattered facts from corpora, while neglecting reasoning complexity, which involves synthesizing interconnected facts into contextually grounded solutions. As shown in Figure 3, they predominantly focus on narrow task categories, such as simple fact retrieval or linear multi-hop reasoning, without systematically capturing the spectrum of challenges encountered in real-world scenarios (Tang & Yang, 2024). For instance, a typical multi-hop question in existing benchmarks might ask, “Who founded Company Kjaer Weis, and in which city was this person born¿‘ This requires only the extraction of several discrete facts and cannot extend to complex scenarios requiring hierarchical reasoning and contextual synthesis. ❷ Second, corpora in existing RAG benchmarks suffer from inconsistent quality and low information density. Many datasets are built on generic sources like Wikipedia or news articles, which lack domain-specific knowledge or explicit logical connections. While some work, like UltraDomain (Qian et al., 2024), has tried to extract domain-specific corpora from textbooks, they often fail to encode implicit hierarchies for real-world applications. This makes it impossible to assess GraphRAG’s core strengths in leveraging domain hierarchies. For example, a corpus with poorly defined conceptual hierarchies or loosely connected entities cannot meaningfully test whether graph-aware retrieval mechanisms improve multi-hop reasoning or preserve contextual coherence during knowledge acquisition. Additionally, the absence of corpora with varying information densities, ranging from tightly structured domain knowledge to loosely organized real-world documents, further restricts the evaluation of graph structures’ scalability and adaptability.
@@
To bridge this gap, we propose GraphRAG-Bench, a comprehensive benchmark designed to evaluate GraphRAG models on deep reasoning. GraphRAG-Bench features ❶ comprehensive corpora with different information density, including tightly structured domain knowledge and loosely organized texts, and ❷ tasks of increasing difficulty, covering fact retrieval, multi-hop reasoning, Contextual Summarize, and creative generation, and ❸ systematic evaluation across the entire pipeline, from graph construction and knowledge retrieval to final generation. Leveraging this novel benchmark, we systematically investigate the conditions when GraphRAG surpasses traditional RAG systems and the underlying reasons for its success, offering guidelines for its practical application.
@@ 2 Preliminary Study
We carefully compare GraphRAG’s pipeline with traditional RAG’s and summarize their characteristics in Figure 1. Generally, RAG retrieves contextually relevant data from a corpus during inference, enabling real-time, domain-specific responses without model retraining. While efficient, its reliance on direct semantic similarity may overlook the broader contextual web of relationships, hierarchies, or implicit logic that binds concepts together. GraphRAG addresses this limitation by expanding the retrieval framework beyond semantic relevance. It structures background knowledge as a graph, where nodes represent entities, events, or themes, and edges define their logical, causal, or associative connections. When processing a query, GraphRAG retrieves not only directly related nodes but also traverses the graph to capture interconnected subgraphs, uncovering latent patterns such as thematic evolution, indirect dependencies, or multi-step reasoning chains. This approach enables the model to synthesize insights from dispersed data points, making it particularly good at tasks demanding complex logical inference. For instance, while RAG might retrieve isolated facts about a topic, GraphRAG could identify related events, causal chains, or thematic clusters, thereby enabling more coherent and comprehensive responses. To sum up, the primary distinction between these two paradigms lies in their handling of contextual depth. RAG excels in scenarios requiring rapid access to discrete information, while GraphRAG emphasizes deep contextual analysis for tasks requiring nuanced understanding of interconnected data. More detailed analysis is included in Appendix I.
@@ 2.2 Current RAG Benchmarks
Existing benchmarks, such as HotpotQA (Yang et al., 2018), MultiHopRAG (Tang & Yang, 2024) and UltraDomain (Qian et al., 2024), were primarily designed to evaluate traditional text-centric RAG frameworks. While these benchmarks have advanced the field, they exhibit critical limitations when applied to assessing GraphRAG.
First, existing benchmarks narrowly focus on testing retrieval difficulty, the ability to locate scattered information from the corpus, while neglecting the equally critical challenge of reasoning difficulty, which involves integrating interconnected concepts/facts by capturing the latent logic. While these benchmarks include “multi-hop“ questions to test a model’s ability in complex reasoning, they do not reflect real-world scenarios demanding complex logical synthesis.
For instance, a typical multi-hop question in existing benchmarks might ask, “Who founded Company Kjaer Weis, and in which city was this person born?” This requires only the extraction of several discrete facts from the corpus. However, real-world problems, such as explaining why Company Kjaer Weis failed in a specific market, demand synthesizing financial reports, competitor analyses, consumer trends, and regulatory changes into a coherent narrative. GraphRAG’s strength lies in mapping these interdependencies (e.g., “Market entry timing → supply chain disruptions → regulatory fines → brand erosion”) through graph traversals. However, current benchmarks lack tasks that explicitly require such synthesis, reducing “multi-hop” queries to sequential fact retrieval within narrow contexts, failing to evaluate how models infer domain-specific hierarchies.
@@
Figure 2: Average number of entities and relations across benchmarks. (Details are in Table 9 in Appendix E)
@@
Second, the corpora used in existing benchmarks suffer from inconsistent quality and low information density. Most datasets are built on generic sources like Wikipedia or news articles, which lack structured domain-specific knowledge or explicit logical connections. A corpus with poorly defined conceptual hierarchies or loosely connected entities cannot meaningfully test whether graph-aware retrieval mechanisms improve multi-hop reasoning or preserve contextual coherence during knowledge integration. While some work, like UltraDomain, has tried to construct domain-specific corpora using textbooks, they often fail to encode implicit hierarchies for real-world applications. As shown in Table LABEL:tab:pre_sta and Table 9 in Appendix E, domain concepts and their hierarchical dependency appear sparsely in the corpus. This sparsity falls far below the threshold of multi-hop reasoning, which makes it impossible to assess GraphRAG’s core strengths in leveraging domain hierarchies. Additionally, the absence of corpora with varying information densities, ranging from tightly structured domain knowledge to loosely organized real-world documents, further restricts the evaluation of graph structures’ scalability.
Third, current benchmarks fall short in evaluating GraphRAG since their evaluation metrics focus solely on the final outputs, answer accuracy or fluency, while treating GraphRAG’s internal processes (graph construction, retrieval, and generation) as black boxes. Such evaluations can hardly measure how graph structures contribute to the retrieval and reasoning processes. To truly assess GraphRAG, a more holistic evaluation is necessary, encompassing the entire pipeline. This includes examining the efficiency of graph construction and the quality of the resulting graph, the structure and relevance of the knowledge retrieved via the graph, and finally, the faithfulness of the generated answer to this graph-derived context. Such a comprehensive view is essential to understand the actual impact and benefits of graph structures within retrieval augmented generation systems. Detailed corpus statistics and analysis for these benchmarks can be found in Section E of the Appendix.
@@ 3 GraphRAG-Bench
In this section, we present GraphRAG-Bench, a novel benchmark specifically designed to assess GraphRAG systems through comprehensive task hierarchies and structured knowledge integration. Specifically, GraphRAG-Bench consists a comprehensive dataset with (i) tasks of increasing difficulty, covering fact retrieval, multi-hop reasoning, Contextual Summarize, and creative generation, and (ii) real-world corpora with different information density, and (iii) a systematic evaluation across the entire pipeline, from graph construction and knowledge retrieval to final generation.